OpenStudy (anonymous):

An online site presented this question, “Would the recent norovirus outbreak deter you from taking a cruise?” Among the 34,696 people who responded, 66% answered “yes”. Use the sample data to construct a 95% confidence interval estimate for the proportion of the population of all people who would respond “yes” to that question. __ < p < __ (round to three decimal places as needed) Does the confidence interval provide a good estimate of the population proportions? (a) No, the sample is a voluntary sample and might not be representative of the population. (b) No, the responses are not independent. (c) Yes, the sample is large enough to provide a good estimate of the population proportion. (d) Yes, all the assumptions for a confidence interval are satisfied.

OpenStudy (anonymous):

Please help me enter this into the Ti84 calc THANKS!!

OpenStudy (anonymous):

We will use the 1-PrpZInt function. We will press the STAT, right arrow key twice. Then, we press ALPHA MATH.

OpenStudy (anonymous):

We will input the appropriate values for x, n and C-level. Then, we will Calculate.

OpenStudy (anonymous):

By the way, take note that p hat = x/n.

OpenStudy (anonymous):

ok, which values are which though...thats the confusing part.

OpenStudy (anonymous):

@violy

OpenStudy (anonymous):

For the p hat, we will take note of the sample proportion given for the people who said yes. can you identify that? For the n, it is the sample size or the number of people who responded.

OpenStudy (anonymous):

yes ok...one sec

OpenStudy (anonymous):

it gave me an error...

OpenStudy (anonymous):

do i put in the 34696 just like that?

OpenStudy (anonymous):

I put x = .66 and n = 34696 then .95 for the c-level and it gave me an error.. :(

OpenStudy (anonymous):

ok so i tried it again and entered just 66 as the x and it gave me (.00144, .00236)..isthat right @violy ?

OpenStudy (anonymous):

Kindly recheck the value of the x.

OpenStudy (anonymous):

The 0.66 is for the p hat. If we use the formula p hat =x/n what will we get for the x?

OpenStudy (anonymous):

.001?

OpenStudy (anonymous):

Oops

OpenStudy (anonymous):

0.66=x/34696. Solve for x.

OpenStudy (anonymous):

Im so clueless, I guess Im not understanding where to get x

OpenStudy (anonymous):

We have the 34696 being divided on right side, we will use the opposite operation of division, which is?

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

That is correct, so we multiply each side by 34696. Can you try it?

OpenStudy (anonymous):

the left side is 22899.36 and the right side is 1203812416

OpenStudy (anonymous):

We will round that.

OpenStudy (anonymous):

ok.....in the question it asks to round to three decimal places as needed...

OpenStudy (anonymous):

@violy